OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (-5,-8) and is perpendicular to 10x-6y=-11

OpenStudy (anonymous):

slope intercept form of an equation of a line is of the form y=mx + b m is the sope b is the intercept first write the equation of the given line in its sope intercept form i.e 10x-6y=-11 is the same as \[y=\frac{ 10 }{ 6x }+ \frac{ 11 }{ 10 }\] from the equation notice the slope of the given line is 10/6 or 5/3. and two lines are perpendicular to one another if the slope of one line is the negative reciprocal to the other so the equation of the perpendicular line is -3/5. the equation of the line perpendicular to 10x - 6y=-11 and passing through (-5,-8) can be found using the point slope form of equation of a line(y-y1=m(x-x1) so we know the slope and the pt it passes through m=-3/5 pt.(-5,-8) y-(-8)=-3/5(x-(-5)) y+8=-3/5x-3 but this is just the point slope form the required is in slope intercept which is y=-3/5x-3-8 y=-3/5x-11 is your answer